OpenStudy (anonymous):

inverse of 9x+3 over x+3?

OpenStudy (amistre64):

swap x and y and solve for y

OpenStudy (amistre64):

9x+3 y = ------ x+3 inverse is: 9y+3 x = ------ y+3 solve for "y" to get it looking "proper"

OpenStudy (anonymous):

you know how to solve for \(y\)?

OpenStudy (anonymous):

multiply both side by y+3?

OpenStudy (anonymous):

that is a good first step, yes \[x(y+3)=9y+3\] next step is to multiply out on the left of the equal sign

OpenStudy (anonymous):

but then you are going to have xy?

OpenStudy (amistre64):

xy is fine, its only a step in the process

OpenStudy (amistre64):

xy + 3x = 9y + 3 get all the y parts to one side and everything else to the other so that we can factor out a common y

OpenStudy (anonymous):

x(y+3)=3(y+1)

OpenStudy (anonymous):

i mean 3(3y+1)

OpenStudy (amistre64):

xy + 3x = 9y + 3 xy - 9y = 3 - 3x y(x - 9) = 3 - 3x and divide off the x-9